Grade 10Trigonometry

in ap the 13th term is 22 and d is -1/2 then first term

saketh
5 Years agoGrade 10

1 Answer

Hari Charan reddy
5 Years ago
we know that, nth term of an AP = a + (n-1)d
so, 13th term = a + (12)d
    i.e. a + 12d = 22
          a +  12(-1/2) = 22
          a – 6 = 22
 Therefore, a(first term) = 22 + 6 = 28.